Find the exact value(s) of x that satisfies the equation 4*3^(2x+4)+124=16000
Then condense everything in your answer to its log.

Answer by Alan3354(69443)

Find the exact value(s) of x that satisfies the equation 4*3^(2x+4)+124=16000
4*3^(2x+4)+124=16000
3^(2x+4)+ 31 = 4000
3^(2x+4) = 3969 = 49*81 = 49*3^4
3^(2x) = 49
3^x = 7
x*log(3) = log(7)
x = log(7)/log(3)